# On Signed Network Coordination Games

### v1

@@ IV-B Stability of Nash equilibria
The network $\mathcal{G}$ in Figure 5 is $(h^{-},h^{+})$-indecomposable for $h^{+}=(3,2,0,2)=-h^{-}$. To see this, first notice that, for any partitiion where $|\mathcal{V}^{+}|=|\mathcal{V}^{-}|=2$, node $3$ satifies either (35) or (36) as a direct consequence of the facts that its out-degree is $w_{3}=3$ and $h_{3}^{-}=h_{3}^{+}=0$. Suppose instead that $|\mathcal{V}^{-}|=1$. A direct check shows that the node in $\mathcal{V}^{-}$ satisfies (36). The case when $|\mathcal{V}^{+}|=1$ is identical.
@@
We will make use of a result in [40] that ensures, for network coordination games on $(h^{-},h^{+})$-indecomposable unsigned networks, the existence of a BR-path from every strategy profile $x$ to a consensus strategy profile that is independent from the specific choice of the vector $h$ satisfying (37). Precisely, [40, Theorem 4(i)] implies the following.
@@
    if, for $y=\pm\bm{1}$ in $\mathcal{X}_{\mathcal{R}}$, the set $\mathcal{N}_{{\mathcal{S}}}^{(y)}$ is globally BR-reachable for the $\mathcal{S}$-restricted SNC game with action profile of players in $\mathcal{R}$ frozen to $y$, then, the subset of Nash equilibria
@@
    if, for $y=\pm\bm{1}$ in $\mathcal{X}_{\mathcal{R}}$, there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(y)}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(y)}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game with action profile of players in $\mathcal{R}$ frozen to $y$, then there exists a non-empty globally BR-stable subset $\overline{\mathcal{N}}\subseteq\tilde{\mathcal{N}}$.
@@
To prove point (ii) of the claim, observe that, if there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(a\bm{1})}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $y=a\bm{1}$, then the BR-path above can be constructed leading to $x^{*}$ such that $x^{*}_{\mathcal{S}}\in\bar{\mathcal{N}}_{{\mathcal{S}}}^{(a\bm{1})}$. Now, notice that assumption (39) implies that, in the strategy profile $x^{*}$ defined above, every player $i$ in ${\mathcal{R}}$ is playing a strict best response $\mathcal{B}_{i}(x^{*}_{-i})=\{a\}$. This, combined with the BR-invariance of $\overline{\mathcal{N}}_{\mathcal{S}}^{(a\bm{1})}$ for the $\mathcal{S}$-restricted SNC game withstrategy profile of players in $\mathcal{R}$ frozen to $y=a\bm{1}$, implies that
@@
Consider a SNC game on the graph in Figure 6 with $h=0$. Observe that =$\mathcal{R}=\{1,\dots,4\}$ and $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}$ are coordinating sets with $w^{\mathcal{R}}|_{\mathcal{R}}-\lvert d\rvert_{\mathcal{R}}=(4,3,3,3)>w^{\mathcal{S}}|_{\mathcal{R}}=(2,2,0,1)$. Furthermore, the graph $\mathcal{G}_{\mathcal{R}}$ is $(h^{+},h^{-})$-indecomposable for $h^{+}=h_{\mathcal{R}}+w^{\mathcal{S}}|_{\mathcal{R}}=(2,2,0,1)$ and $h^{-}=h_{\mathcal{R}}-w^{\mathcal{S}}|_{\mathcal{R}}=-h^{+}$. This can be proved following the same reasoning as in Example 7 (notice that $\mathcal{G}_{\mathcal{R}}$ coincides with the graph in Figure 7). Then, according to Corollary 3, the set of Nash equilibria where the players in ${\mathcal{R}}$ are at consensus is globally BR-reachable. Notice that the set of Nash equilibria
@@
Consider the graph in Figure 4. As observed in Example 6 subset =$\mathcal{R}=\{1,\dots,4\}$ is such that the graph $\mathcal{G}_{\mathcal{R}}$ is structurally balanced. Indeed, if we consider the gauge transformation $[\sigma]$ with $\sigma=[1,-1,-1,1,1,1]$, we obtain that $W^{[\sigma]}=[\sigma]W[\sigma]\geq 0$. Furthermore, for $h_{i}=0$, we have $w^{\mathcal{R}}|_{\mathcal{R}}=(4,3,4,3)>w^{\mathcal{S}}|_{\mathcal{R}}=(1,1,1,1)$. Since $\mathcal{G}_{\mathcal{S}}$ is a undirected (no links), this implies that, for $\tau=(1,-1,-1,1)$, there exists a polarized Nash equilibrium, that is, a Nash equilibrium where $x_{i}^{*}=\tau_{i}$ for all $i$ in $\mathcal{R}$. For instance, the strategy profile $x^{*}=(1,-1,-1,1,-1,1)$ is a Nash equilibrium of the game.
Finally, since $\tilde{\mathcal{G}}_{\mathcal{R}}$ with $\tilde{W}_{\mathcal{R}\mathcal{R}}=\lvert W_{\mathcal{R}\mathcal{R}}\rvert$ is $(h^{+},h^{-})$-indecomposable for $h^{+}=\bm{1}$ and $h^{-}=-\bm{1}$, the set of Nash equilibria that are polarized in the structurally balanced component admits a globally BR-stable subset. We remark that these results can be obtained without knowing the partitions $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$.
@@ Appendix A Proof of Lemma 1
First, observe that $x^{*}$ in $\mathcal{X}$ is a Nash equilibrium for the $[\sigma]$-transformed game if and only if
$x_{i}^{*}\left(\sum\nolimits_{j}W^{[\sigma]}_{ij}x^{*}_{j}+h^{[\sigma]}_{i}\right)\geq 0\,,\qquad\forall i\in\mathcal{V}\,.$
@@
$\begin{array}[]{rcl}\!\!\!\displaystyle x_{i}^{*}\left(\sum\nolimits_{j}W^{[\sigma]}_{ij}x^{*}_{j}+h^{[\sigma]}_{i}\right)&\!\!\!\!=\!\!\!\!&x_{i}^{*}\left(\displaystyle\sum\nolimits_{j}\sigma_{i}W_{ij}\sigma_{j}x^{*}_{j}+\sigma_{i}h_{i}\right)\\[10.0pt] &\!\!\!\!=\!\!\!\!&\displaystyle\sigma_{i}x_{i}^{*}\left(\sum\nolimits_{j}W_{ij}\sigma_{j}x^{*}_{j}+h_{i}\right)\\[10.0pt] &\!\!\!\!=\!\!\!\!&\displaystyle([\sigma]x^{*})_{i}\left(\sum\nolimits_{j}W_{ij}([\sigma]x^{*})_{j}+h_{i}\right).\end{array}$
Therefore, $x^{*}$ is a Nash equilibrium for the $[\sigma]$-transformed game if and only if $[\sigma]x^{*}$ is a Nash equilibrium for the SNC game with binary actions on $\mathcal{G}$ with external field $h$.
@@ Appendix B Proof of Lemma 2
[Uncaptioned image] Martina Vanelli received the B.Sc., M.S. (cum laude), and Ph.D. (with honors) degrees in Applied Mathematics from Politecnico di Torino, in 2017, 2019, and 2024, respectively. She is currently a postdoctoral fellow at the Institute for Information and Communication Technologies, Electronics and Applied Mathematics (ICTEAM), Université catholique de Louvain. From October 2018 to March 2019, she was a visiting student at Technion, Israel Institute of Technology. Her research interests include identification, analysis, and control of multi-agent systems, with application to social and economic networks, and power markets.
@@
[Uncaptioned image] Giacomo Como is a Professor at the Department of Mathematical Sciences, Politecnico di Torino, Italy, and a Senior Lecturer at the Automatic Control Department of Lund University, Sweden. He received the B.Sc., M.S., and Ph.D. degrees in Applied Mathematics from Politecnico di Torino, in 2002, 2004, and 2008, respectively. He was a Visiting Assistant in Research at Yale University in 2006–2007 and a Postdoctoral Associate at the Laboratory for Information and Decision Systems, Massachusetts Institute of Technology, from 2008 to 2011. He currently serves as Senior Editor of the IEEE Transactions on Control of Network Systems and Associate Editor of Automatica. . He has been serving as Associate Editor of the IEEE Transactions on Network Science and Engineering and of the IEEE Transactions on Control of Network Systems and as the chair of the IEEE-CSS Technical Committee on Networks and Communications. He was the IPC chair of the IFAC Workshop NecSys’15 and a semi-plenary speaker at the International Symposium MTNS’16. He is recipient of the 2015 George S.  Axelby Outstanding Paper Award. His research interests are in dynamics, information, and control in network systems with applications to cyber-physical systems, infrastructure networks, and social and economic networks.

### v2

@@ IV-B Stability of Nash equilibria
The network $\mathcal{G}$ in Figure 5 is $(h^{-},h^{+})$-indecomposable for $h^{+}=(3,2,0,2)=-h^{-}$. To see this, first notice that, for any partition where $|\mathcal{V}^{+}|=|\mathcal{V}^{-}|=2$, node $3$ satifies either (36) or (37) as a direct consequence of the facts that its out-degree is $w_{3}=3$ and $h_{3}^{-}=h_{3}^{+}=0$. Suppose instead that $|\mathcal{V}^{-}|=1$. A direct check shows that the node in $\mathcal{V}^{-}$ satisfies (37). The case when $|\mathcal{V}^{+}|=1$ is identical.
@@
We will make use of a result in [41] that ensures, for network coordination games on $(h^{-},h^{+})$-indecomposable unsigned networks, the existence of a BR-path from every strategy profile $x$ to a consensus strategy profile that is independent from the specific choice of the vector $h$ satisfying (38). Precisely, [41, Theorem 4(i)] implies the following.
@@
    if, for every $a$ in {±$\{\pm 1\}$, the set $\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$ is globally BR-reachable for the $\mathcal{S}$-restricted SNC game with action profile of players in $\mathcal{R}$ frozen to $a\bm{1}$, then, the subset of Nash equilibria
@@
    if, for every $a$ in {±$\{\pm 1\}$, there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(a\bm{1})}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game with action profile of players in $\mathcal{R}$ frozen to $a\bm{1}$, then there exists a non-empty globally BR-stable subset $\overline{\mathcal{N}}\subseteq\tilde{\mathcal{N}}$.
@@
To prove point (ii) of the claim, let $a$ in {±$\{\pm 1\}$ be as above. Since there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(a\bm{1})}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $a\bm{1}$, then the BR-path above can be constructed leading to $x^{*}$ such that $x^{*}_{\mathcal{S}}\in\bar{\mathcal{N}}_{{\mathcal{S}}}^{(a\bm{1})}$. Now, notice that assumption (40) implies that, in the strategy profile $x^{*}$ defined above, every player $i$ in ${\mathcal{R}}$ is playing a strict best response $\mathcal{B}_{i}(x^{*}_{-i})=\{a\}$. This, combined with the BR-invariance of $\overline{\mathcal{N}}_{\mathcal{S}}^{(a\bm{1})}$ for the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $a\bm{1}$, implies that
@@
Consider a SNC game on the graph in Figure 6 with $h=0$. Observe that =$\mathcal{R}=\{1,\dots,4\}$ and $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}$ are coordinating sets with $w^{\mathcal{R}}|_{\mathcal{R}}-\lvert h\rvert_{\mathcal{R}}=(4,3,3,3)>w^{\mathcal{S}}|_{\mathcal{R}}=(2,2,0,1)$. Furthermore, the graph $\mathcal{G}_{\mathcal{R}}$ is $(h^{+},h^{-})$-indecomposable for $h^{+}=h_{\mathcal{R}}+w^{\mathcal{S}}|_{\mathcal{R}}=(2,2,0,1)$ and $h^{-}=h_{\mathcal{R}}-w^{\mathcal{S}}|_{\mathcal{R}}=-h^{+}$. This can be proved following the same reasoning as in Example 7 (notice that $\mathcal{G}_{\mathcal{R}}$ coincides with the graph in Figure 7). Then, according to Corollary 3, the set of Nash equilibria where the players in ${\mathcal{R}}$ are at consensus is globally BR-reachable. Notice that the set of Nash equilibria
@@
Consider the graph in Figure 4. As previously observed, the subset =$\mathcal{R}=\{1,\dots,4\}$ is such that the graph $\mathcal{G}_{\mathcal{R}}$ is structurally balanced. Indeed, if we consider the gauge transformation $[\sigma]$ with $\sigma=[1,-1,-1,1,1,1]$, we obtain that $W^{[\sigma]}=[\sigma]W[\sigma]\geq 0$. Furthermore, recall that $\mathcal{G}_{\mathcal{S}}$ is undirected.
For $h=0$, we have that $w^{\mathcal{R}}|_{\mathcal{R}}=(4,3,4,3)>w^{\mathcal{S}}|_{\mathcal{R}}=(1,1,1,1)$, which implies that (40) holds true. Furthermore, we find that $\mathcal{G}_{\mathcal{R}}$ with $\tilde{W}_{\mathcal{R}\mathcal{R}}=\lvert W_{\mathcal{R}\mathcal{R}}\rvert$ is $(h^{+},h^{-})$-indecomposable for $h^{+}=h+w^{\mathcal{S}}|_{\mathcal{R}}=\bm{1}$ and $h^{-}=h-w^{\mathcal{S}}|_{\mathcal{R}}=-\bm{1}$. Then, Corollary 4 applies and the set of Nash equilibria that are polarized in the structurally balanced component admits a globally BR-stable subset.
@@ Appendix A Proof of Lemma 1
First, observe that $a$ in $\mathcal{A}$ is a best response for a player $i$ in $\mathcal{V}$ to a strategy profile $x_{-i}$ in $\mathcal{X}_{-i}$ for the $[\sigma]$-transformed game, if and only if
$a\left(\sum\nolimits_{j}W^{[\sigma]}_{ij}x_{j}+h^{[\sigma]}_{i}\right)\geq 0\,,\qquad\forall i\in\mathcal{V}\,.$
@@
$\begin{array}[]{rcl}\!\!\!\displaystyle a\left(\sum\nolimits_{j}W^{[\sigma]}_{ij}x_{j}+h^{[\sigma]}_{i}\right)&\!\!\!\!=\!\!\!\!&a\left(\displaystyle\sum\nolimits_{j}\sigma_{i}W_{ij}\sigma_{j}x_{j}+\sigma_{i}h_{i}\right)\\[10.0pt] &\!\!\!\!=\!\!\!\!&\displaystyle\sigma_{i}a\left(\sum\nolimits_{j}W_{ij}\sigma_{j}x_{j}+h_{i}\right)\\[10.0pt] &\!\!\!\!=\!\!\!\!&\displaystyle\sigma_{i}a\left(\sum\nolimits_{j}W_{ij}([\sigma]x)_{j}+h_{i}\right)\,.\end{array}$
Therefore, $a\in\mathcal{B}_{i}^{[\sigma]}(x_{-i})$ if and only if $\sigma_{i}a\in\mathcal{B}_{i}(([\sigma]x)_{-i})$, thus proving (17).
It then follows from (17) that $x^{*}$ is a Nash equilibrium for the $[\sigma]$-transformed game, i.e., $x_{i}^{*}\in\mathcal{B}_{i}^{[\sigma]}(x^{*}_{-i})$, if and only if $\sigma_{i}x_{i}^{*}\in\mathcal{B}_{i}([\sigma]x^{*})_{-i})$, i.e., $[\sigma]x^{*}$ is a Nash equilibrium for the SNC game with binary actions on $\mathcal{G}$ with external field $h$. This proves (18).
@@ Appendix B Proof of Lemma 2
[Uncaptioned image] Martina Vanelli received the B.Sc., M.S. (cum laude), and Ph.D. (with honors) degrees in Applied Mathematics from Politecnico di Torino, in 2017, 2019, and 2024, respectively. She is currently a postdoctoral fellow at the Institute for Information and Communication Technologies, Electronics and Applied Mathematics (ICTEAM), Université catholique de Louvain. From October 2018 to March 2019, she was a visiting student at Technion, Israel Institute of Technology. Her research interests include analysis and control of multi-agent and network systems, primarily with application to social and economic networks, and power markets.
@@
[Uncaptioned image] Giacomo Como is a Professor at the Department of Mathematical Sciences, Politecnico di Torino, Italy, and a Senior Lecturer at the Automatic Control Department of Lund University, Sweden. He received the B.Sc., M.S., and Ph.D. degrees in Applied Mathematics from Politecnico di Torino, in 2002, 2004, and 2008, respectively. He was a Visiting Assistant in Research at Yale University in 2006–2007 and a Postdoctoral Associate at the Laboratory for Information and Decision Systems, Massachusetts Institute of Technology, from 2008 to 2011. He currently serves as Senior Editor of the IEEE Transactions on Control of Network Systems and Associate Editor of Automatica. He has been serving as Associate Editor of the IEEE Transactions on Network Science and Engineering and of the IEEE Transactions on Control of Network Systems and as the chair of the IEEE-CSS Technical Committee on Networks and Communications. He was the IPC chair of the IFAC Workshop NecSys’15 and a semi-plenary speaker at the International Symposium MTNS’16. He is recipient of the 2015 George S. Axelby Outstanding Paper Award. His research interests are in dynamics, information, and control in network systems with applications to cyber-physical systems, infrastructure networks, and social and economic networks.